OpenStudy (anonymous):

find and sketch the domain of the function in 3D: f(x,y) = sqrt(x*y)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

so far all I got is x*y cannt be less than 0

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

xy < 0 means that either x < 0 or y < 0

jimthompson5910 (jim_thompson5910):

if both x < 0 AND y < 0, then xy > 0

jimthompson5910 (jim_thompson5910):

so you just have to make sure that either case 1) both x and y are positive (or zero) OR case 2) both x and y are negative (or zero)

OpenStudy (anonymous):

i need to sketch it :/

OpenStudy (anonymous):

@jim_thompson5910 do you know how to sketch it in 3D?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

im having trouble understanding how to sketch functions in 3D

jimthompson5910 (jim_thompson5910):

without software, I think it's going to be a real pain to sketch in 3D

jimthompson5910 (jim_thompson5910):

hopefully you can use a computer tool?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i found what it's supposed to look like online but it's such a wierd shape lol

jimthompson5910 (jim_thompson5910):

well I guess you could think of either x or y to be constant while the other is a variable

jimthompson5910 (jim_thompson5910):

so for instance, let y be fixed to y = 2 sqrt(x*y) = sqrt(x*2) = sqrt(2x)

OpenStudy (anonymous):